OpenStudy (anonymous):

plaese help me! a physical education teacher was getting ready for gym class one day. he measured the circumference of the basketball to be 9 inches when its filled with air. he then went on to pose the following questions to his classes: A) how much material is needed to make four basketball? (3.14 for pie)

OpenStudy (anonymous):

b) how much air will the four balls hold combined? ( 3.14 for pie)

OpenStudy (anonymous):

c) if the basketballs are stored four to a box in cubic boxes whose edges are 15 incges long, what percent of the box is not filled by balls?

jimthompson5910 (jim_thompson5910):

part a), you're finding the surface area of the ball or sphere

jimthompson5910 (jim_thompson5910):

do you know the formula for the surface area of a sphere?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok what is it

jimthompson5910 (jim_thompson5910):

well first off, you'll need to find the radius before you can find the surface area

OpenStudy (anonymous):

4(3.14)r squared

jimthompson5910 (jim_thompson5910):

C = 2*pi*r 9 = 2*3.14*r solve for r

jimthompson5910 (jim_thompson5910):

good, keep that formula in mind for later

OpenStudy (anonymous):

okay the problem u just out up is for A right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

r=4.5

jimthompson5910 (jim_thompson5910):

9 = 2*3.14*r 9/(2*3.14) = r 1.43312 = r r = 1.43312

jimthompson5910 (jim_thompson5910):

now you use this in 4(3.14)r squared to find the surface area

OpenStudy (anonymous):

i got 25.796 for surface area

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

that was for B right??

jimthompson5910 (jim_thompson5910):

now multiply that answer by 4 why? because that answer of 25.796 is the surface area of one basketball but you want to make 4 basketballs

jimthompson5910 (jim_thompson5910):

no all still on part a

OpenStudy (anonymous):

103.184

jimthompson5910 (jim_thompson5910):

that's your final answer to part a

jimthompson5910 (jim_thompson5910):

that's how much material you need to make the four basketballs

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

for part B do i use Volume?

jimthompson5910 (jim_thompson5910):

yes you use the formula V = (4/3)*pi*r^3

jimthompson5910 (jim_thompson5910):

that's to get the volume of one ball

OpenStudy (anonymous):

i got 49.2916 for 4 balls

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

getting the same pretty much

jimthompson5910 (jim_thompson5910):

so the volume of 4 balls is roughly 49.2916 cubic inches

OpenStudy (anonymous):

yes :) i don't get what i have to do for part C??

jimthompson5910 (jim_thompson5910):

what is the volume of the box?

OpenStudy (anonymous):

i don't know

jimthompson5910 (jim_thompson5910):

V = s^3 V = 15^3 V = 3375 cubic inches

OpenStudy (anonymous):

okay i got it :) i have one last question :)

jimthompson5910 (jim_thompson5910):

well we're not done with C, do you see how to get the answer for C?

jimthompson5910 (jim_thompson5910):

using what I gave you

OpenStudy (anonymous):

what do i do next ??

jimthompson5910 (jim_thompson5910):

the volume of the four balls was 49.2916

jimthompson5910 (jim_thompson5910):

the volume of the box is 3375

jimthompson5910 (jim_thompson5910):

so the volume inside the box that is NOT taken up by the 4 balls is 3375 - 49.2916 = 3325.7084

jimthompson5910 (jim_thompson5910):

therefore, the percent of the box NOT filled by the balls is 3325.7084/3375 = 0.985395 = 98.5395%

jimthompson5910 (jim_thompson5910):

so you have a lot of room in this box

OpenStudy (anonymous):

got it :)

OpenStudy (anonymous):

the school would like to sane money ans storage space by buying smaller boxes to store the balls. what is the smallest size for the box where it will still hold four balls and each edge will be a whole number/

OpenStudy (anonymous):

last question :)

jimthompson5910 (jim_thompson5910):

ok say the box had side lengths of 1 inch this would give it a volume of 1^3 = 1 cubic inches that's too small since we need something that will hold 49.2916 cubic inches

jimthompson5910 (jim_thompson5910):

2^3 = 8 ... still too small 3^3 = 27 ... again too small 4^3 = 64 ... not too small

jimthompson5910 (jim_thompson5910):

so the smallest box has side lengths of 4 inches

OpenStudy (anonymous):

thank you soo much :)

jimthompson5910 (jim_thompson5910):

sure thing